# Insights from leptohadronic modelling of the brightest blazar flare

### v1

@@ 2.5 Assumed systematic uncertainties
From the comparison of measured magnitudes by SMARTS and Steward for the same object for the same night, we found a typical value of $\approx 0.08$ for the relative uncertainty of the SED (see Appendix A). We add in quadrature 8% relative systematic uncertainty to the statistical uncertainty of all IR, optical and UV data points. Moreover, both Fermi-LAT and Swift-XRT have some uncertainty in the estimations of their effective areas resulting in the systematic uncertainty of the measured SEDs. We assume $3\%$ relative systematic uncertainty, which can be expected for both Fermi-LAT [^10] https://fermi.gsfc.nasa.gov/ssc/data/analysis/scitools/Aeff_Systematics.html and Swift-XRT (Moretti et al., 2006, sect. 5), adding it in quadrature to all the uncertainties of all the analysed SEDs (including IR, optical, and UV data points).
@@ 3 Modelling
$x_{s}$ [cm] Blob launching distance from the SMBHm $f$m
$x_{f}$ [cm] Blob final distance from the SMBHm $f$m
@@
$R_{\mathrm{AD,\,in}}\,[R_{g}]$ Inner radius of the accretion disc $3$
@@
$t^{o}_{\mathrm{obs}}$ [s] Total time of observations in the observer’s rest framem $d$m
$t_{\mathrm{tot}}$ [s] Total time of the blob motion in the host galaxy rest framem $d$m, see Eq. (27)
$t^{o}_{\mathrm{shift}}$ [s] Time delay of the observations w.r.t. the blob launchm $d$m
@@
Table 1: Definitions of variables mentioned throughout the text. Primed quantities $[^{\prime}]$ denote variables measured in the blob rest frame. Unprimed variables are either invariant or in the host galaxy rest frame. A superscript o denotes variables measured in the Earth rest frame at the present cosmological epoch. A letter $d$ denotes variables whose values are derived from the other variables, and a letter $f$ denotes variables that are free during the optimization. A superscript m denotes variables used in the model with a single “moving blob”. A superscript h denotes the average value of those reported by Ghisellini et al. (2010); Nalewajko et al. (2019); Titarchuk et al. (2020).
We assume that the observed multi-wavelength spectral-energy distribution (SED $=E^{2}dN/dE=\nu F_{\nu}=EF(E)$) can be modelled as the radiation emitted by a population of relativistic electrons (including positrons which we do not distinguish from electrons unless explicitly stated otherwise) trapped in the turbulent magnetic field of a spherical plasmoid (blob) located in the jet at a distance $x$ from the supermassive black hole (SMBH) where the plasma flows with bulk Lorentz factor $\Gamma$ with a small angle $\theta_{\mathrm{view}}$ to the LoS (see Fig. 1).
For modelling the observable SED, we use the time-dependent program for the simulation of radiative processes AM3 (Klinger et al., 2023, 2024), although other solutions are also available (see, e.g., the review of programs simulating radiative processes by Nigro & Tramacere, 2022). AM3 is a versatile program for time-dependent modelling of various radiative processes (both leptonic and hadronic) including the synchrotron and inverse-Compton (IC) processes (for leptonic models), and the photopion and Bethe-Heitler (BH) processes (for leptohadronic models) with the subsequent cooling of the charged particles. Throughout the paper, we use the variables with their notations explained in Table 1. We assume the fiducial model of Gilmore et al. (2012) for the calculation of the absorption of $\gamma$ rays on the extragalactic background light (EBL) on their way from 3C 454.3 to Earth and the flat $\Lambda$CDM Universe with the Hubble parameter $H_{0}=67.74$ km/(s Mpc) and the normalized matter density $\Omega_{m}=0.3075$ (Planck Collaboration et al., 2016) resulting in the luminosity distance to 3C 454.3 $d_{L}=1.74\times 10^{28}$ cm given its redshift $z=0.859$ (Jackson & Browne, 1991).
Using the time-dependent solver AM3, we model the observed SEDs of the MJD period 55517–55524 in two approaches: 1) assuming that the radiation comes from a region located at a particular distance $x$ from the SMBH which does not change throughout the modelled period, i.e. the emission region is attributed to a stationary feature in the jet – hereafter the “standing-feature” model; 2) alternatively, assuming that the radiation comes from a single blob which moves along with the bulk motion of the jet plasma – hereafter the “moving-blob” model. We perform parameter optimization utilizing iMINUIT (Dembinski et al., 2021; James & Roos, 1975) and emcee (Foreman-Mackey et al., 2013) as well as develop routines to estimate values of some model parameters as described in the following section.
@@ 3.1 Description of the “standing-feature” model
$\beta=\sqrt{1-\Gamma^{-2}}$, and $\theta_{\mathrm{view}}=0.4^{\circ}$, is the angle between the jet direction and the LoS (the assumption is based on the radio data by, e.g., Kovalev et al., 2020; Pushkarev et al., 2017; Jorstad et al., 2013). We assume that the injection of electrons stops after $t^{\prime}_{\mathrm{sim}}$ and – since their energy-loss timescale is for most relevant energies much less than $t^{\prime}_{\mathrm{sim}}$ – their emission stops shortly after that.
Following Rodrigues et al. (2024), we model all radiative processes in the blob rest frame having transformed external photon fields into this frame [^12] Though, an alternative sequence of frame transformations is possible (e.g., Khangulyan et al., 2018).. The obtained number densities $n^{\prime}(E^{\prime})$, differential per volume and per energy, are transformed into the observer frame SED $\nu F_{\nu}$ as (see, e.g., Jiménez-Fernández & van Eerten (2021), eq. 1)
@@ 1. Accretion disc:
    Direct radiation of the accretion disc (AD) in the model of Shakura & Sunyaev (1976) is calculated according to Ghisellini & Tavecchio (2009, sect. 3.1) and Ehlert et al. (2024, appendix C1). On various estimates of the 3C 454.3 AD luminosity in literature, see Table 2. We adopt $L_{\mathrm{AD}}=3\times 10^{46}$ erg/s which is approximately the median value of those given in Table 2.
@@ 2. Broad line region:
    Radiation intercepted and reprocessed by the BLR modelled as the superposition of the H I Ly $\alpha$ and He II Ly $\alpha$ lines and thermal continuum according to the examples in the AM3 documentation (Klinger et al., 2023, 2024; Rodrigues et al., 2024) with the continuum re-emitting $C_{\mathrm{BLR}}=0.01$ of the AD luminosity, where $C_{\mathrm{BLR}}$ is the BLR continuum reprocessing factor. In addition, the BLR lines reprocess $\xi_{\mathrm{BLR}}=0.10$ of the AD luminosity where $\xi_{\mathrm{BLR}}$ is the BLR covering factor (Baldwin & Netzer, 1978). We calculate the BLR outer radius using the standard relation (Kaspi et al., 2007; Ghisellini & Tavecchio, 2009)
@@ 3. Dust Torus:
    Radiation intercepted and reprocessed by the dust torus (DT) region is modelled as a thermal continuum with temperature $T_{\mathrm{DT}}=370$ K and covering factor $C_{\mathrm{DT}}=0.30$ according to the examples in the AM3 documentation (Klinger et al., 2023, 2024; Rodrigues et al., 2024). The DT outer radius is assumed to follow (Błażejowski et al., 2000; Sikora et al., 2002; Ghisellini & Tavecchio, 2009)
@@ 3.1.2 Escape timescales
For a leptonic setup, the only particles involved in the simulation are photons and electrons. The timescale for photons escaping from the blob is (see, e.g., Jiménez-Fernández & van Eerten, 2021, eq. 2):
@@ 3.1.5 Location of the $\gamma$-ray emitting region
In this subsection, we obtain an analytical estimate of $x$ to use as an initial guess before the iMINUIT optimization. The distance $x$ of the region emitting the observed radiation from the SMBH establishes the energy at which the high-energy $\gamma$ rays will be absorbed by the photons of the BLR photon field.
@@
[…]
@@
$\Gamma_{\gamma}$ $2.67\pm 0.07$
$\chi^{2}_{\mathrm{red}}$ $0.77$
@@
$\Gamma_{\gamma}$ (before the break) $2.1$
$\Gamma_{\gamma}$ (after the break) $2.6$
$\chi^{2}_{\mathrm{red}}$ $2.83$
@@
$\gamma_{1}$ $1.92$
$\gamma_{b}$ $0.51$
$\chi^{2}_{\mathrm{red}}$ $0.89$
@@ 3.1.6 Magnetic field and parameters of the electron injection spectrum
For the Fermi-LAT data, it is clear that they cannot be described by a single power law (see also, e.g., Rajput et al., 2019; Sahakyan, 2021). We failed to obtain a good fit with a broken power law (BPL) with $\chi^{2}_{\mathrm{red}}=2.83$. On the contrary, a fit with a subexponential cutoff [^15] The Fermi-LAT Collaboration uses the term “superexponential cutoff” https://fermi.gsfc.nasa.gov/ssc/data/analysis/scitools/source_models.html, but in our case $\gamma_{b}<1$, so we use the term “subexponential” instead.
$\mathrm{SED}(E)=E^{2}N_{0}\left(\frac{E}{E_{0}}\right)^{-\gamma_{1}}\exp{\left(-\left(\frac{E}{E_{\mathrm{cut}}}\right)^{\gamma_{b}}\right)}$ (21)
describes the Fermi-LAT range quite well ($\chi^{2}_{\mathrm{red}}=0.89$). We interpret this as evidence for the curvature of the spectrum reported previously by Abdo et al. (2011); Pacciani et al. (2014); Sahakyan (2021). Indeed, in the 4FGL catalogue itself, the spectral model for 3C 454.3 is chosen to have a subexponential cutoff (Abdollahi et al., 2020) providing the best description of the eight-year data. Besides, Stern & Poutanen (2014) from the analysis of almost five years of Fermi-LAT data, found that the SED of 3C 454.3 is best described by a biquadratic function
@@
$E^{\prime\mathrm{max}}_{e}=\frac{1+z}{\delta}E^{o}_{\gamma_{\mathrm{max}}},$ (23)
so $E^{\prime\mathrm{max}}_{e}\approx 1.5~{}\times~{}10^{9}$ eV for $E^{o}_{\gamma_{\mathrm{max}}}\approx 4~{}\times~{}10^{10}$ eV.
From the broken power-law fit to the Fermi-LAT data in Fig. 4, we can assume the power-law slope of the cooled electron spectrum $dN_{e}/dE_{e}\propto E^{-p_{e}}$ to be (Blumenthal & Gould, 1970; Finke & Dermer, 2010; Ghisellini et al., 2020)
@@
Assuming that the UV photons at the energy $E^{\mathrm{o}}_{\mathrm{UV}}\approx 4$ eV are exclusively produced by the cooled electrons at the cutoff $E^{\prime\mathrm{max}}_{e}$, we can estimate the root-mean-square magnetic field strength as (e.g., Jiménez-Fernández & van Eerten, 2021, eq. 55)
$B^{\prime}\approx\frac{4\pi m_{e}^{3}c^{5}E^{\mathrm{o}}_{\mathrm{UV}}}{hq_{e}\delta\left(E^{\prime\mathrm{max}}_{e}\right)^{2}},$ (26)
@@
Finally, as initial guess we assume $E^{\prime\mathrm{min}}_{e}=10^{8}$ eV and $L^{\prime}_{e}=2\times 10^{43}$ erg/s.
With the model parameters’ values from our initial guess, we run the AM3 program with the total simulation time according to Eq. (4). The injected and cooled electron SED at the end of the simulation are shown in Fig. 6 along with the fit of the functional form of Eq. (21) to the high-energy part of the cooled electron SED resulting in excellent agreement with the superexponential cutoff with $\gamma_{b}=3.7$. The superexponential cutoff of the cooled electron SED in Fig. 6 results in the subexponential cutoff of the $\gamma$-ray SED with $\gamma_{b}=0.51$ (see Fig. 5), which is in agreement with the expected value from Lefa et al. (2012) for a case in between the synchrotron ($\gamma_{b}=3.7/(3.7+4)=0.48$) and Planckian ($\gamma_{b}=3.7/(3.7+2)=0.65$) photons serving as the target photon fields (Lefa et al., 2012, table 1) for the IC in the Thomson regime.

### v2

@@ 2.5 Assumed systematic uncertainties
From the comparison of measured magnitudes by SMARTS and Steward for the same object for the same night, we found a typical value of $\approx 0.08$ for the relative uncertainty of the SED (see Appendix A). We add in quadrature 8% relative systematic uncertainty to the statistical uncertainty of all IR, optical and UV data points. Moreover, both Fermi-LAT and Swift-XRT have some uncertainty in the estimations of their effective areas, which results in a systematic uncertainty in the measured SEDs. We assume $3\%$ relative systematic uncertainty, which can be expected for both Fermi-LAT [^10] https://fermi.gsfc.nasa.gov/ssc/data/analysis/scitools/Aeff_Systematics.html and Swift-XRT (Moretti et al., 2006, sect. 5), adding it in quadrature to all the uncertainties of all the analysed SEDs (including IR, optical, and UV data points; Tables 1, 2, and 3 show values before adding any systematic uncertainty).
@@ 3 Leptonic models
$x_{s}$ [cm] Blob launching distance from the SMBHm ffm
$x_{f}$ [cm] Blob final distance from the SMBHm ffm
@@
$R_{\mathrm{AD,\,in}}\,[R_{g}]$ Inner radius of the accretion disc 33
@@
$t^{o}_{\mathrm{obs}}$ [s] Total time of observations in the observer’s rest framem ddm
$t_{\mathrm{tot}}$ [s] Total time of the blob motion in the host galaxy rest framem ddm, see Eq. (26)
$t^{o}_{\mathrm{shift}}$ [s] Time delay of the observations w.r.t. the blob launchm ddm
@@
Table 4: Definitions of variables mentioned throughout the text. Primed quantities $^{\prime}$] denote variables measured in the blob rest frame. Unprimed variables are either invariant or in the host galaxy rest frame. A superscript o denotes variables measured in the Earth rest frame at the present cosmological epoch. A letter $d$ denotes variables whose values are derived from the other variables, and a letter $f$ denotes variables that are free during the optimisation. A superscript m denotes variables used in the model with a single “moving blob”. A superscript h denotes the average value of those reported by Ghisellini et al. (2010); Nalewajko et al. (2019); Titarchuk et al. (2020).
We assume that the observed multi-wavelength spectral-energy distribution (SED $E^{2}dN/dE=\nu F_{\nu}=EF(E)$) can be modelled as the radiation emitted by a population of relativistic electrons (including positrons which we do not distinguish from electrons unless explicitly stated otherwise) trapped in the turbulent magnetic field of a spherical plasmoid (blob) located in the jet at a distance $x$ from the supermassive black hole (SMBH) where the plasma flows with bulk Lorentz factor $\Gamma$ with a small angle $\theta_{\mathrm{view}}$ to the LoS (see Fig. 1).
[…]
Using the time-dependent solver AM3, we model the observed SEDs of the MJD period 55517–55524 in two approaches: 1) assuming that the radiation comes from a region located at a particular distance $x$ from the SMBH which does not change throughout the modelled period, i.e. the emission region is attributed to a stationary feature in the jet – hereafter the “standing-feature” model; 2) alternatively, assuming that the radiation comes from a single blob which moves along with the bulk motion of the jet plasma – hereafter the “moving-blob” model. We perform parameter optimisation utilizing iMINUIT (Dembinski et al., 2021; James & Roos, 1975) and emcee (Foreman-Mackey et al., 2013) as well as develop routines to estimate values of some model parameters as described in the following section.
@@ 3.1 Description of the “standing-feature” model
$\beta=\sqrt{1-\Gamma^{-2}}$, and $\theta_{\mathrm{view}}=0.4^{\circ}$, is the angle between the jet direction and the LoS (the assumption is based on the radio data by, e.g., Kovalev et al., 2020; Pushkarev et al., 2017; Jorstad et al., 2013). We assume that the injection of electrons stops after $t^{\prime}_{\mathrm{sim}}$ and — since their energy-loss timescale is for most relevant energies much less than $t^{\prime}_{\mathrm{sim}}$ — their emission stops shortly after that. We verify that in the optical/UV and $\gamma$-ray range, the steady-state SED is reached after around half of $t^{\prime}_{\mathrm{sim}}$ (or, in other words, after around three light-crossing (dynamical) timescales), i.e., after that time, the optical/UV and $\gamma$-ray SED do not change. However, the X-ray SED somewhat depends on $t^{\prime}_{\mathrm{sim}}$ because the timescale for the SSC process is comparable to the simulation timescale. We find that the photon flux in the energy range between $0.2$ keV and $10$ keV varies by $\pm 15$% after changing the simulation time by a factor of $2$, which is less than or comparable to the relative uncertainty of the model parameters. Thus, we assume that by the end of the AM3 simulation, we obtain a steady-state model SED, which is fitted to the corresponding daily-averaged quasi-simultaneous observed SED.
Following Rodrigues et al. (2024), we model all radiative processes in the blob rest frame having transformed external photon fields into this frame [^12] Though, an alternative sequence of frame transformations is possible (e.g., Khangulyan et al., 2018).. The obtained number densities $n^{\prime}(E^{\prime})$, differential per volume and per energy, are transformed into the observer frame SED $\nu F_{\nu}$ as (see, e.g., Dermer & Menon (2009), sect. 5.5.1)
@@ 1. Accretion disc:
    Direct radiation of the accretion disc (AD) in the model of Shakura & Sunyaev (1976) is calculated according to Ghisellini & Tavecchio (2009, sect. 3.1) and Ehlert et al. (2025, appendix C1). On various estimates of the 3C 454.3 AD luminosity in literature, see Table 5. We adopt $L_{\mathrm{AD}}=3\times 10^{46}$ erg/s, which is approximately the median value of those given in Table 5.
@@ 2. Broad line region:
    Radiation intercepted and reprocessed by the BLR modelled as the superposition of the H I Ly $\alpha$ (10.2 eV), He II Ly $\alpha$ (40.8 eV) lines, and thermal continuum, according to the examples in the AM3 documentation (Klinger et al., 2023, 2024; Rodrigues et al., 2024), with the continuum re-emitting $C_{\mathrm{BLR}}=0.01$ (e.g., Blandford & Levinson, 1995) of the AD luminosity, where $C_{\mathrm{BLR}}$ is the BLR continuum reprocessing factor. In addition, the BLR H I Ly lines reprocess $\xi_{\mathrm{BLR}}=0.10$ of the AD luminosity, and He II Ly — $0.5\xi_{\mathrm{BLR}}$ of the AD luminosity, where $\xi_{\mathrm{BLR}}$ is the BLR covering factor (Baldwin & Netzer, 1978; D’Elia et al., 2003). The lines are assumed to be emitted at the same $R_{\mathrm{BLR}}$ distance from the SMBH and to have the relative Gaussian widths of 55% of their central values. It is the H I Ly line which is the most luminous while the continuum and other lines have significantly lower contribution to the external photon field density and have much higher uncertainty of their relative luminosity (Finke, 2016, appendix, and table 5). For simplicity, our BLR model follows Rodrigues et al. (2024), who followed Murase et al. (2014), where they claim that their results depend weakly on the ratio of helium-over-hydrogen line luminosities, which is also observationally uncertain (e.g., Zheng & Fang, 1998). We check that a factor-of-a-few change of the relative contribution of He II Ly or BLR continuum to the external photon field density does not change the observable SED substantially, which confirms the claims of Murase et al. (2014). We calculate the BLR outer radius using the standard relation (Kaspi et al., 2007; Ghisellini & Tavecchio, 2009)
@@ 3. Dust Torus:
    Radiation intercepted and reprocessed by the dust torus (DT) region is modelled as a thermal continuum with temperature $T_{\mathrm{DT}}=370$ K [^14] Typical values of $T_{\mathrm{DT}}$ are in between 1002000100-2000 K (Cleary et al., 2007; Malmrose et al., 2011). In our best-fitting “standing-feature” model the DT photon field is subdominant with respect to the BLR photon field, so a factor-of-a-few change of $T_{\mathrm{DT}}$ does not cause any substantial change in the observable SED. and covering factor $C_{\mathrm{DT}}=0.30$ according to the examples in the AM3 documentation (Klinger et al., 2023, 2024; Rodrigues et al., 2024). The DT outer radius is assumed to follow (Błażejowski et al., 2000; Sikora et al., 2002; Ghisellini & Tavecchio, 2009)
@@ 3.1.2 Escape timescales
For a leptonic setup, the only particles involved in the simulation are photons and electrons. The timescale for photons escaping from the blob is
@@
We verify that the confined electrons after $t^{\prime}_{\mathrm{sim}}$, when the injection is stopped, do not produce any substantial signal which could affect the following days. This is expected because, in the energy range of electrons responsible for observable electromagnetic emission, they have an energy-loss timescale much shorter than the simulation timescale.
@@ 3.1.5 Location of the $\gamma$-ray emitting region
In this subsection, we obtain an analytical estimate of $x$ to use as an initial guess before the iMINUIT optimisation. The distance $x$ of the region emitting the observed radiation from the SMBH establishes the energy at which the high-energy $\gamma$ rays will be absorbed by the photons of the BLR photon field.
@@
$\tau_{\gamma\gamma}(E^{o}_{\gamma},x)=\int_{x}^{+\infty}r^{\prime}_{\gamma\gamma}(E^{\prime}_{\gamma},y)dy,$ (13)
@@
$\Gamma_{\gamma}$ $2.52\pm 0.03$
$\chi^{2}_{\mathrm{red}}$ $0.39$
@@
$\Gamma_{\gamma}$ (before the break) $2.076\pm 0.014$
$\Gamma_{\gamma}$ (after the break) $2.60\pm 0.04$
$E_{\mathrm{break}}$ [GeV] $1.00\pm 0.09$
$\chi^{2}_{\mathrm{red}}$ $2.63$
@@
$\gamma_{1}$ $1.896\pm 0.026$
$\gamma_{b}$ $0.49\pm 0.04$
$E_{\mathrm{cut}}$ [GeV] $1.85\pm 0.16$
$\chi^{2}_{\mathrm{red}}$ $0.80$
@@ 3.1.6 Magnetic field and parameters of the electron injection spectrum
For the Fermi-LAT data, it is clear that they cannot be described by a single power law (see also, e.g., Rajput et al., 2019; Sahakyan, 2021). We failed to obtain a good fit with a broken power law (BPL) with $\chi^{2}_{\mathrm{red}}=2.63$. On the contrary, a fit with a subexponential cutoff (compare with eq. 1),
$\mathrm{SED}(E)=E^{2}\frac{dN}{dE}=E^{2}N_{0}\left(\frac{E}{E_{0}}\right)^{-\gamma_{1}}\exp{\left(-\left(\frac{E}{E_{\mathrm{cut}}}\right)^{\gamma_{b}}\right)},$ (19)
describes the Fermi-LAT range quite well ($\chi^{2}_{\mathrm{red}}=0.80$). We interpret this as evidence for the curvature of the spectrum reported previously by Abdo et al. (2011); Pacciani et al. (2014); Sahakyan (2021). Indeed, in the 4FGL catalogue itself, the spectral model for 3C 454.3 is chosen to have a subexponential cutoff (Abdollahi et al., 2020) providing the best description of the eight-year data. Besides, Stern & Poutanen (2014) from the analysis of almost five years of Fermi-LAT data, found that the SED of 3C 454.3 is best described by a biquadratic function
@@
$E^{\prime\mathrm{max}}_{e}=\frac{1+z}{\delta}E^{o}_{\gamma_{\mathrm{max}}}$ (21)
in the Klein-Nishina regime, or, in the Thomson regime (e.g., Dermer & Menon, 2009, eq. 6.52),
$E^{\prime\mathrm{max}}_{e}=m_{e}c^{2}\sqrt{\frac{3E^{o}_{\gamma_{\mathrm{max}}}(1+z)}{4\delta\langle\epsilon_{\mathrm{BLR}}^{\prime}\rangle}},$ (22)
where $\langle\epsilon_{\mathrm{BLR}}^{\prime}\rangle$ is the average energy of BLR photons in the blob rest frame, which can be estimated with Eqs. (15), and (16) for $\epsilon\approx 20$ eV corresponding to the average between H I Ly $\alpha$ and He II Ly $\alpha$ lines. For the observed maximum photon energy $E^{o}_{\gamma_{\mathrm{max}}}\approx 4~\times~10^{10}$ eV, we obtain $E^{\prime\mathrm{max}}_{e}\approx 1.5~\times~10^{9}$ eV according to Eq. (21) and a similar value of $E^{\prime\mathrm{max}}_{e}\approx 1.7~\times~10^{9}$ eV according to Eq. (22). To be more conservative, we choose the smaller value as our initial guess [^16] We also check that the electron inverse Compton scattering off the He II Ly $\alpha$ photons enters the Klein-Nishina regime while the scattering off the H I Ly $\alpha$ photons is still in the Thomson regime..
From the broken power-law fit to the Fermi-LAT data in Fig. 4, we can assume the power-law slope of the cooled electron spectrum $dN_{e}/dE_{e}\propto E^{-p_{e}}$ to be (Blumenthal & Gould, 1970; Rybicki & Lightman, 1979)
@@
Assuming that the UV photons at the energy $E^{\mathrm{o}}_{\mathrm{UV}}\approx 4$ eV are exclusively produced by the cooled electrons at the cutoff $E^{\prime\mathrm{max}}_{e}$, we can estimate the root-mean-square magnetic field strength as (e.g., Rybicki & Lightman, 1979, eq. 6.17c)
$B^{\prime}\approx\frac{4\pi m_{e}^{3}c^{5}E^{\mathrm{o}}_{\mathrm{UV}}(1+z)}{3hq_{e}\delta\left(E^{\prime\mathrm{max}}_{e}\right)^{2}},$ (25)
@@
For the initial guess of the minimum electron injection energy $E^{\prime\mathrm{min}}_{e}$, we assume that the observed by Fermi-LAT peak of the $\gamma$-ray SED at around $E^{o}_{\gamma_{\mathrm{peak}}}\approx 2\times 10^{8}$ eV is due to the break in the cooled electron spectrum happening at $E^{\prime\mathrm{min}}_{e}$, which, given the good approximation of the IC process by the Thomson formula at lower energies, can be estimated with the same Eq. (22) as for the maximum electron injection energy, just by substituting $E^{o}_{\gamma_{\mathrm{max}}}$ with $E^{o}_{\gamma_{\mathrm{peak}}}$. This gives us the initial guess $E^{\prime\mathrm{min}}_{e}\approx 10^{8}$ eV.
Finally, for the last parameter — electron injection luminosity — we set $L^{\prime}_{e}=2\times 10^{43}$ erg/s as our initial guess, just comparing the model and observed SEDs by eye, trying several values of $L^{\prime}_{e}$.
With the model parameters’ values from our initial guess, we run the AM3 program with the total simulation time according to Eq. (2). The injected and cooled electron SEDs at the end of the simulation are shown in Fig. 6 along with the fit of the functional form of Eq. (19) to the high-energy part of the cooled electron SED, resulting in excellent agreement with the superexponential cutoff with $\gamma_{b}=3.7$. The superexponential cutoff of the cooled electron SED in Fig. 6 results in the subexponential cutoff of the $\gamma$-ray SED with $\gamma_{b}=0.49$ (see Fig. 5 and Table 6), which is in agreement with the expected value from Lefa et al. (2012) for a case in between the synchrotron ($\gamma_{b}=3.7/(3.7+4)=0.48$) and Planckian ($\gamma_{b}=3.7/(3.7+2)=0.65$) photons serving as the target photon fields (Lefa et al., 2012, table 1) for the IC in the Thomson regime.